OpenStudy (anonymous):

Please help. i narrowed it down from what i have done. Solve: 2 3/4+X+3 1/2=10. solve

OpenStudy (lasttccasey):

is that supposed to be 2.75+x+3.5=10?

OpenStudy (anonymous):

yes

OpenStudy (lasttccasey):

Okay so subtract the 2.75 and 3.5 to the right side leaving x on the left, x = 10 - 2.75 - 3.5 x = 3.75

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

= 3 3/4 as a fraction.

OpenStudy (lasttccasey):

You're welcome and yes, 3.75 = 3 3/4 or 15/4